OpenStudy (anonymous):

The probability that Shania is on time for school is 1/2 . Find the probability that Shania arrives on time for school for the next 5 days. Express your answer as a percent, to the nearest tenth of a percent.

OpenStudy (blacksteel):

Shania arriving on time each day is an independent probability (ie, the probability of arriving on time on Tuesday is not affected by whether she got there on time on Monday). Therefore, the probability of arriving on time 5 days in a row is 0.5*0.5*0.5*0.5*0.5, or (1/2)^5 = 0.03125, or 3.125% Round as necessary

OpenStudy (callisto):

P(Shania arrives on time for school for the next 5 days) = P(arrived on time on first day) x P(arrived on time on second day) x P(arrived on time on third day) x P(arrived on time on fourth day) x P(arrived on time on fifth day) = (1/2) x (1/2) x (1/2) x (1/2) x (1/2) = (1/2)^5 = ?

OpenStudy (anonymous):

1/32

OpenStudy (callisto):

And the decimal number is?

OpenStudy (anonymous):

0,03125

OpenStudy (anonymous):

oh. I understend. Thank you very much

OpenStudy (callisto):

Welcome :) remember to convert it into percentage by multiply 100%. The round to the nearest 10th, that is the digit after the ''.'' :)

OpenStudy (anonymous):

thank you, Callisto:)

OpenStudy (callisto):

Welcome :)

OpenStudy (callisto):

I can't figure out why it wasn't 1/21 for your first question I answered. Maybe I need to think twice about it!

OpenStudy (anonymous):

I don't understand too